Find the 5th term and the sum of the first 5 terms of the geometric sequence 2, -1/2,1/8,....

Determine the first terms of the terms of the sequence of {(-1/n)power to n} and give their corresponding sum.
1
Expert's answer
2020-12-01T20:43:56-0500

q=bn+1/b=-0.5/2=-0.25

b5=b1qn-1=2x(-0.25)5-1=2/256=1/125

"S_n=\frac{b_1(1-q^n)}{q-1}=\frac{2(1-(-0.25)^5}{(-0.25)-1} =\frac{2(1+1/1024)}{-5/4}=\frac{8(1+1/1024)}{-5}=-(8/5 +1/128) =- 1.61"

First terms of (-1/n)n are: -1, 1/4, -1/27

"q=\frac{1/4}{-1}=-1/4"

"S_3=\frac{-1(1-(-1/4)^3)}{(-0.25-1)}=\frac{-65/64}{-5/4}=\frac{65x4}{64x5}=\frac{13}{16}"